OpenStudy (anonymous):

find the locus of a point p(x,y) which moves so that its distance from the point (1,0) is half its distance from the line x=-1

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

use the distance formula

OpenStudy (anonymous):

\[\sqrt{(x-1)^2+(y-0)^2}=\frac{1}{2}\sqrt{(x+1)^2+(y-y)^2}\]

OpenStudy (anonymous):

no 2x^2-4x+2+2y^2=x^2+2x+1

OpenStudy (dumbcow):

i get an ellipse, gonna double check it \[\frac{6(x-\frac{5}{3})^{2}}{19} + \frac{8y^{2}}{19} = 1\]

OpenStudy (anonymous):

x^2+2y^2-2x+1=0

OpenStudy (dumbcow):

nevermind im wrong

OpenStudy (anonymous):

Thanks alot though!... And Purplec16 Thanks! got that too i jus wanted to double check.

OpenStudy (dumbcow):

hmm purple you have to square the 1/2 as well when you square both sides 4y^2 +4x^2 -8x +4 = x^2 +2x +1

OpenStudy (dumbcow):

found my mistake, its half of an ellipse \[\frac{9(x-\frac{5}{3})^{2}}{16} + \frac{3y^{2}}{4} = 1\] for x<= 5/3